If you made an investment of 2000 today about how long would it take for it to double in value if interest rate on investment is %10?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

(2*2000)=2000(1.1)^n

2=(1.1)^n

Taking log on both sides;

log 2=n*log 1.1

Hence n=log 2/log 1.1

which is equal to

=7.27 years(Approx).